At 25c,0.752mol of Ar gas occupied 0.274 L.what is the pressure,in atm, of the Area gas?
1
Expert's answer
2018-05-14T08:14:03-0400
pV = nRT
p = nRT/V = 0.752mol* 8.31J/mol*K * (25+273)K / 0,000274 m^3 = 6.79*10^6 Pa.